A bomb at rest explodes into 2 parts of mass m1 and m2. If the momentums of 2 parts be p1 and p2. then kinetic energy is.........? Explain it briefly .

Grade:11

1 Answers

downey zer
6 years ago
we know p=mv and KE=(1/2)mv 
=(1/2)(mv)2/(m)
=p2/2m
therfore total kE =p12/2m1 +  p22/2m2 
but p1 = -p2 (since the bomb was at rest initially and no external force has been applied on the system therefore momentum is conserved)
total kE= p12/2m1 + p12/2m2 = p22/2m1 + p22/2m2
please approve the answer if it helped